# DivCon: Divide and Conquer for Progressive Text-to-Image Generation

### v1

@@ Abstract
Diffusion-driven text-to-image (T2I) generation has achieved remarkable advancements. To further improve T2I models’ capability in numerical and spatial reasoning, the layout is employed as an intermedium to bridge large language models and layout-based diffusion models. However, these methods still struggle with generating images from textural prompts with multiple objects and complicated spatial relationships. To tackle this challenge, we introduce a divide-and-conquer approach which decouples the T2I generation task into simple subtasks. Our approach divides the layout prediction stage into numerical & spatial reasoning and bounding box prediction. Then, the layout-to-image generation stage is conducted in an iterative manner to reconstruct objects from easy ones to difficult ones. We conduct experiments on the HRS and NSR-1K benchmarks and our approach outperforms previous state-of-the-art models with notable margins. In addition, visual results demonstrate that our approach significantly improves the controllability and consistency in generating multiple objects from complex textural prompts.
Keywords:
Text-to-Image Generation Large Language Models Diffusion Models
Figure 1: Layouts and images generated by our DivCon. DivCon enhances the ability of text-to-image diffusion models to understand complex numerical and spatial relationships in the text.
1 Introduction
@@
Large-scale text-to-image generation (T2I) models demonstrate exceptional zero-shot capacity in efficiently generating small quantities and varieties of objects [30, 7, 3, 21, 29, 31]. However, these methods still suffer from limited capability in generating images from text prompts containing specified object counts, varying sizes, rich details, and complicated spatial relationships [2, 7, 14, 8, 35]. Recently, combining Large Language Models (LLMs) and layout-to-image diffusion models has significantly improved T2I models’ ability in numerical and spatial reasoning [15, 26]. LLMs have strong reasoning and visual planning ability to predict objects’ layouts from the text prompt, which provides more informative conditions for diffusion models to produce higher-quality images.
@@
Despite the significant advancement, existing methods still suffer from limited accuracy when the text prompt contains multiple objects and complicated spatial relationships [2, 26]. Intuitively, the difficulty of synthesizing objects with different characteristics varies a lot. Specifically, LLMs encounter great challenges when conducting simultaneous reasoning and visual planning from complicated text prompts (e.g. "five mice are scurrying around a wooden table, each nibbling on one of the five ripe apples that are laid out on top."). Besides, layout-to-image diffusion models exhibit varying levels of generative proficiency for objects with different sizes, shapes, and details. However, previous methods generate all objects simultaneously without taking their differences into consideration. As a result, objects with higher difficulties cannot be well generated in the images (e.g. toaster and pizza, Fig. 2).
@@
To address this issue, we propose DivCon, a training-free approach that applies the divide-and-conquer strategy to both layout prediction and image generation phases. Our approach divides complicated tasks into multiple subtasks according to their difficulties and then address these subtasks in an iterative manner. During layout prediction, the number of objects and their spatial relationships are first reasoned from the text prompt. Then, visual planning is performed to predict object bounding boxes. During layout-to-image generation, the diffusion model is first executed to process the layout. Then, the consistency between the synthetic objects and the text prompt is calculated. Next, low-fidelity objects’ layouts are highlighted and fed to the diffusion model to encourage the network to focus on objects with higher difficulties. Extensive experiments have demonstrated that DivCon has significantly improved the quality of generated and achieves state-of-the-art performance on the HRS[2] and NSR-1K[15] benchmark datasets.
@@
  • We propose DivCon, a divide-and-conquer approach for layout-based text-to-image generation by dividing complicated tasks into simple subtasks.
  • To obtain accurate layouts from the text prompt, we divide the layout prediction stage into numerical & spatial reasoning and bounding box planning.
  • To generate high-fidelity images from the layout, we divide layout-to-image generation into two iterations to synthesize objects with different levels of difficulty.
@@
Figure 2: A comparison of generation difficulty between "toaster" and "pizza". Current methods tend to generate pizza but fail to generate enough toasters or completely miss the category.
2 Related work
In this section, we review three main categories of text-to-image generation methods: text-to-image generation (T2I) models, layout-conditioned T2I models, and text-to-layout-to-image generation models (Fig. 3).
2.0.1 Text-to-Image (T2I) Generation Models
The evolution of large-scale T2I synthesis demonstrates a significant leap forward in the realm of generative models, enabling the creation of detailed and coherent images directly from textual descriptions [3, 17, 29, 21, 31, 38]. Milestone methods in this area are Generative Adversarial Networks (GANs) [21, 32], Variational Autoencoders (VAEs) [6, 12, 29, 38], and diffusion models [3, 19, 25, 28, 31]. GANs often struggle with training stability and mode collapse, whereas VAEs may sometimes produce overly smooth or blurry outputs due to their reconstruction loss. Recently, diffusion models are highly valued for their robustness and ability to generate detailed and realistic images. Particularly, Stable Diffusion[30], which utilizes open vocabulary encoder[27] for encoding text to embeddings as conditional inputs, has demonstrated a remarkable capability for open-world image synthesis, pushing the boundaries of generative models. Despite promising results, a text prompt is usually difficult to provide fine-grained control of the generated images.
Figure 3: Three main categories of text-to-image generation methods.
2.0.2 Layout-conditioned T2I Models
To improve the controlability of T2I models, prompts like layouts as employed as additional conditions for T2I models. Existing layout-conditioned generation models can be categorized into three branches. One branch of methods trains external networks on layout-images pairs [22, 39, 25, 1, 37]. For instance, GLIGEN incorporates a trainable gated self-attention layer that transforms bounding box information into conditional input for Stable Diffusion. Another branch of methods, including attention-refocusing [26], layout-predictor [34], direct-diffusion[24], layout-guidance[9] and BoxDiff [36], directly manipulate attention maps to optimize the sampling process. These methods generate images with multiple objects through a single forward pass of the diffusion model, resulting in limited fidelity for objects with complicated appearances or spatial relationships. To remedy this, the last branch of methods including Mixture-of-Diffusion[5], MultiDiffusion [4], and LLM-grounded Diffusion [23] perform denoising on each object (i.e., bounding box) separately before fusing the results. However, these methods introduce much higher computational costs. Moreover, the requirements of additional layout prompts hinder their practicality.
2.0.3 Text-to-Layout-to-Image Generation Models
Current LLMs have demonstrated strong numerical and spatial reasoning abilities in visual planning and layout generation. Inspired by this, several efforts are made to integrate LLMs to generate layouts from texts, and then utilize layout-conditioned diffusion models for image generation. Attention-refocusing [26] and LLM-grounded Diffusion [23] explore in-context learning of GPT-4 for layout prediction. LayoutGPT [15] searches similar samples from their database as input examples for the LLMs. LLM-layout-generator [10] leverages Chain-of-Thought (CoT) prompting [33] to improve the layout prediction accuracy. However, LLMs are expected to simultaneously perform numerical & spatial reasoning and visual planning of bounding boxes in these methods. Contrary to these methods, we propose a divide-and-conquer approach to conduct reasoning and visual planning separately, which improves the accuracy of layout predictions and substantially advances the prompt fidelity of generated images. In addition, during layout-to-image generation, the objects are synthesized in an iterative manner according to their difficulties, which achieves not only superior fidelity but also significant computational cost savings.
@@ 2.0.3 Text-to-Layout-to-Image Generation Models
Figure 4: The proposed DivCon framework. The layout prediction stage is divided into numerical & spatial reasoning and bounding box planning. The layout-to-image generation is divided into two iterations to generate objects with different difficulty levels separately.
3 Method
3.1 Overview
@@
As illustrated in Fig. 4, our framework comprises two stages. Given a text prompt, LLMs first predict layouts from it. Then, the resultant layout is employed as an additional condition to achieve text-to-image generation. More specifically, each stage is further divided into two steps to handle simple subtasks.
@@ 3.2 Text-to-Layout Prediction
Intuitively, to predict the layout from a text prompt, the number & spatial relations of the objects are first required to be reasoned, and then bounding boxes are generated to describe the coordinates & sizes of the objects. However, previous methods commonly conduct these two steps simultaneously without considering their inherent correlation, thereby hindering them from understanding complicated text prompts well. To remedy this, we introduce the idea of divide-and-conquer to improve the accuracy of predicted layouts. Specifically, leveraging the state-of-the-art large language model GPT-4 [], numerical and spatial reasoning is first performed, and then bounding boxes are predicted.
3.2.1 Numerical and Spatial Reasoning
Given a text prompt $C$ describing an image, an LLM is adopted to predict a set of strings $O=\{s_{j}|j=1,2,...,n\}$. If the prompt primarily focused on counting relationships among objects, each string $s_{j}$ includes both the name and quantity information of each object category $j$. If the prompt mainly describes spatial relationships, each string $s_{j}$ will contain the object name and its spatial position in the image. This method allows LLMs to conclude complex numerical and spatial relationships.
3.2.2 Bounding Box Prediction
After numerical and spatial reasoning, both the meta prompt $C$ and the prediction result $O$ are fed to the LLM again to predict a list of name-box pairs $P=\{p_{i}|i=1,2,...,n\}$. Each pair $p_{i}=\{o_{i}:(l_{i},t_{i},r_{i},b_{i})|i=1,2,...,n\}$ consists of the object name $o_{i}$ and its bounding box’s coordinates $(l_{i},t_{i},r_{i},b_{i})$. The in-context exemplars include a process of correcting the features of generated bounding boxes to ensure the quality of the output layouts.
@@ 3.3 Layout-to-Image Generation
After layout prediction, the resultant layout is employed as an additional condition to generate the image. As objects in the layout have different levels of difficulty to synthesize, a divide-and-conquer approach is adopted to reconstruct objects from easy to hard in an iterative manner, as shown in Fig. 4. First, the layout is fed to the diffusion model to synthesize an image. Then, the consistency between the synthetic objects and the text prompt is calculated. Afterward, objects with low fidelity are highlighted and passed to the diffusion model for a second time, encouraging the diffusion model to focus on these difficult objects to achieve higher quality.
3.3.1 First-Round Generation
@@
We employ a layout-to-image diffusion model with sampling optimization [26] as the generative model to generate an image conditioned on the text prompt and the predicted layout prompt. During the denoising process, random noise $z_{T}$ is fed to the diffusion model to gradually denoise it to an image $z_{0}$ after $T$ iterations.
3.3.2 Consistency Evaluation
@@
With the generated image $x$, each object is cropped using the bounding boxes in the layout, which are denoted as $x_{i}\in\mathbb{R}^{4\times h\times w}$, where $h\times w$ represents the height and width of the bounding box. Then, the CLIP similarity $S_{x_{i},o_{i}}\in(-1,1)$ is employed to evaluate the consistency between the cropped patch and the object name $o_{i}$ in the text prompt. The higher the similarity is, the higher the fidelity of the corresponding object is. Consequently, the objects with similarities higher than a specified threshold can be considered as easy samples that are "good enough”. In contrast, other objects with low similarities are hard samples that are more difficult to reconstruct (e.g. "sheep" and "horse" in Fig. 5). Next, a binary mask $M$ is calculated with bounding boxes of easy objects being set to 1 and the rest regions being set to 0.
Figure 5: Comparison of easy and hard samples in a generated image.
3.3.3 Second-Round Refinement
@@
After consistency evaluation, different objects are of various levels of fidelity. Our objective is to preserve the “good enough” objects while reproducing the remaining ones to ease the difficulty of diffusion model. To this end, we retrieve the latent representation during the denoising process of the first-round generation. Specifically, $z_{0.7T}$ and $z_{0.3T}$ are first retrieved. Then, the binary mask $M$ is adopted to crop high-fidelity regions in $z_{0.7T}$ and $z_{0.3T}$. Next, the starting noise $z^{\prime}_{T}$ for the second-round diffusion is obtained as:
$z^{\prime}_{T}=z_{0.7T}\circ M+\mathcal{N}(\mu,\sigma^{2})\circ(1-M)$ (1)
where $\circ$ is element-wise multiplication and $\mathcal{N}(\mu,\sigma^{2})$ is random noise.
The second-round denoising process consists of three stages. First, the input noise $z^{\prime}_{T}$ is denoised for $0.3T$ iterations to obtain $z^{\prime}_{0.3T}$. Second, the high-fidelity regions in $z_{0.7T}$ are cropped and used to replace the corresponding regions in $z^{\prime}_{0.3T}$ and fed to the diffusion model for anthoer $0.4T$ denoising iterations. Note that, the text prompt and the full layout prompt are employed as conditions in these two stages. Moreover, the regions marked by $M$ is frozen to preserve these high-fidelity regions. During optimization, Cross-Attention Refocusing (CAR) and Self-Attention Refocusing (SAR) losses [26] are employed to update only regions of low fidelity, encouraging the model to focus on these hadr samples. Finally, all regions are optimized for the last $0.3T$ iterations to smoothly blend the high-fidelity regions with other regions.
4 Experiments
@@
In this section, we provide an evaluation of our method and compare it with state-of-the-art T2I models. Ablation experiments are conducted to demonstrate the effectiveness of our divide-and-conquer approach in layout prediction and layout-to-image generation stages.

### v2

@@ Abstract
Diffusion-driven text-to-image (T2I) generation has achieved remarkable advancements in recent years. To further improve T2I models’ capability in numerical and spatial reasoning, layout is employed as an intermedium to bridge large language models and layout-based diffusion models. However, these methods still struggle with generating images from prompts with multiple objects and complicated spatial relationships. To tackle this challenge, we introduce a divide-and-conquer approach which decouples the generation task into multiple subtasks. First, the layout prediction stage is divided into numerical & spatial reasoning and bounding box prediction to better parse the text prompt to obtain accurate composition information. Second, the layout-to-image generation stage is divided into two steps to synthesize objects from easy ones to difficult ones. Experiments are conducted on the HRS and NSR-1K benchmarks and our method outperforms previous approaches with notable margins. In addition, visual results and user study demonstrate that our approach significantly improves the perceptual quality, especially when generating multiple objects from complex textural prompts.
Codehttps://github.com/DivCon-gen/DivCon
Introduction
@@
Large-scale text-to-image generation (T2I) models demonstrate exceptional zero-shot capacity in efficiently generating small quantities and varieties of objects (Rombach et al. 2022; Chefer et al. 2023; Balaji et al. 2022; Kang et al. 2023; Ramesh et al. 2021; Saharia et al. 2022). However, these methods still suffer limited capability in generating images from text prompts containing specified object counts, varying sizes, rich details, and complicated spatial relationships (Bakr et al. 2023; Chefer et al. 2023; Feng et al. 2023a; Chen, Laina, and Vedaldi 2023a; Xiao et al. 2023). Recently, combining large language models (LLMs) and layout-to-image diffusion models has significantly improved T2I models’ capability in numerical and spatial reasoning (Feng et al. 2023b; Phung, Ge, and Huang 2024). These methods leverage the strong reasoning and visual planning ability of LLMs to predict objects’ layouts from the text prompt, which provides additional informative conditions for diffusion models to produce higher-quality images.
@@
Despite recent significant advancements, existing methods still suffer limited accuracy when the text prompt contains multiple objects and complicated spatial relationships (Bakr et al. 2023; Phung, Ge, and Huang 2024). On the one hand, LLMs encounter great challenges when conducting simultaneous reasoning and visual planning from complicated text prompts (e.g. “five mice are scurrying around a wooden table, each nibbling on one of the five ripe apples that are laid out on top”). On the other hand, since the difficulty of synthesizing objects with different characteristics varies a lot, layout-to-image diffusion models exhibit varying levels of generative proficiency for objects with different sizes, shapes, and details. However, previous methods commonly conduct reasoning and visual planning together while generating all objects simultaneously without taking their differences into consideration. As a result, objects with higher difficulties cannot be well synthesized in the images (e.g. toaster and pizza in Fig. 2).
Figure 1: Layouts and images generated by our DivCon. DivCon enhances the capability of text-to-image diffusion models to understand complex numerical and spatial relationships in the text.
@@
To address this issue, we propose DivCon, a training-free approach that applies the divide-and-conquer strategy to both layout prediction and image generation stages. During layout prediction, the number of objects and their spatial relationships are first reasoned from the text prompt. Then, visual planning is conducted to predict the bounding boxes of the objects. By decoupling layout prediction into reasoning and visual planning, more accurate composition information can be obtained. During layout-to-image generation, the diffusion model is first executed to process the layout to generate all objects. Then, the consistency between the resultant objects and the text prompt is calculated. With objects of low consistency being highlighted, the diffusion model is executed for another time to encourage the network to focus more on these difficult objects. By diving layout-to-image generation into two steps, objects with different levels of difficulties can be synthesized progressively for higher quality. Extensive experiments have demonstrated that our DivCon has significantly improved the quality of generated images and achieves state-of-the-art performance on the HRS(Bakr et al. 2023) and NSR-1K(Feng et al. 2023b) benchmark datasets.
@@
  • We propose DivCon, a divide-and-conquer approach for layout-based text-to-image generation by dividing the complicated task into multiple subtasks.
  • To obtain accurate composition information from the text prompt, we divide the layout prediction stage into numerical & spatial reasoning and bounding box planning.
  • To generate high-fidelity images from the layout, we divide layout-to-image generation into two steps to synthesize objects with different levels of difficulty.
@@
Figure 2: A comparison of generation difficulty between “toaster” and “pizza”.
Figure 3: Comparison of easy and hard samples in a generated image.
@@
Figure 4: The proposed DivCon framework. The layout prediction stage (stage 1) is divided into numerical & spatial reasoning and bounding box planning. The layout-to-image generation stage (stage 2) is divided into two steps to generate objects with different difficulty levels separately.
Related Work
In this section, we review three main categories of text-to-image generation methods: text-to-image generation (T2I) models, layout-conditioned T2I models, and text-to-layout-to-image generation models.
2.1 Text-to-Image (T2I) Generation Models
The evolution of large-scale T2I synthesis demonstrates a significant leap forward in the realm of generative models, enabling the creation of detailed and coherent images directly from textual descriptions (Balaji et al. 2022; Gafni et al. 2022; Ramesh et al. 2021; Kang et al. 2023; Saharia et al. 2022; Yu et al. 2022). Milestone methods in this area are Generative Adversarial Networks (GANs) (Kang et al. 2023; Sauer et al. 2023), Variational Autoencoders (VAEs) (Chang et al. 2023; Ding et al. 2022; Ramesh et al. 2021; Yu et al. 2022), and diffusion models (Balaji et al. 2022; Ho, Jain, and Abbeel 2020; Nichol et al. 2022; Ramesh et al. 2022; Saharia et al. 2022). GANs often struggle with training stability and mode collapse, whereas VAEs may sometimes produce overly smooth or blurry outputs due to their reconstruction loss. Recently, diffusion models are highly valued for their robustness and ability to generate detailed and realistic images. Particularly, Stable Diffusion (Rombach et al. 2022), which utilizes text embeddings extracted by CLIP (Radford et al. 2021) as conditions, has demonstrated remarkable performance for open-world image synthesis. Despite promising results, a text prompt is usually difficult to provide fine-grained control of the generated images.
2.2 Layout-conditioned T2I Models
To improve the controlability of T2I models, prompts like layouts are employed as additional conditions for T2I models. Existing layout-conditioned generation models can be categorized into three branches. One branch of methods train external networks on layout-images pairs (Li et al. 2023b; Zhang, Rao, and Agrawala 2023; Nichol et al. 2022; Avrahami et al. 2023; Yang et al. 2022). For instance, GLIGEN incorporates a trainable gated self-attention layer that transforms bounding boxes into conditional input for Stable Diffusion. Another branch of methods, including attention-refocusing (Phung, Ge, and Huang 2024), layout-predictor (Wu et al. 2023), direct-diffusion(Ma et al. 2023), layout-guidance(Chen, Laina, and Vedaldi 2023b) and BoxDiff (Xie et al. 2023), directly manipulate attention maps to optimize the sampling process of the diffusion model. Since these methods generate all objects through a single forward pass of the diffusion model, objects with spatial relationships usually suffer limited quality. To remedy this, the last branch of methods (including Mixture-of-Diffusion(Barbero Jiménez 2023), MultiDiffusion (Bar-Tal et al. 2023), and LLM-grounded Diffusion (Lian et al. 2023)) first synthesize each object (i.e., bounding box) separately and then compose them to produce the final result. Nevertheless, the computational cost of these methods increases with the number of objects such that the computational overhead is considerable for a text prompt containing plenty of objects. In addition, the relationship between objects is disrupted and cannot be well maintained in the generated images.
2.3 Text-to-Layout-to-Image Generation Models
Current LLMs have demonstrated strong numerical and spatial reasoning capabilities in visual planning and layout generation. Inspired by this, several efforts are made to integrate LLMs to generate layouts from texts, and then utilize layout-conditioned diffusion models for image generation. Attention-refocusing (Phung, Ge, and Huang 2024) and LLM-grounded Diffusion (Lian et al. 2023) explore in-context learning of GPT-4 for layout prediction. LayoutGPT (Feng et al. 2023b) retrieves similar samples from a database as input examples for LLMs. LLM-layout-generator (Chen et al. 2023) leverages Chain-of-Thought (CoT) prompting (Wei et al. 2022) to improve the layout prediction accuracy. However, LLMs are expected to simultaneously perform numerical & spatial reasoning and visual planning of bounding boxes in these methods. As a result, inferior accuracy is produced especially for complicated text prompts. Contrary to these methods, we propose a divide-and-conquer approach to conduct reasoning and layout prediction separately, which improves the accuracy of layout predictions and substantially advances the prompt fidelity of generated images. In addition, during layout-to-image generation, the objects are synthesized progressively according to their difficulties, which achieves not only superior fidelity but also significant computational cost savings.
Methods Numerical Layout Pre. Rec. F1 Acc. Image Pre. Rec. F1 Acc. Spatial Layout Image Acc. Acc.
Stable Diffusion 2.1 - - - - 73.97 53.93 62.38 16.20 - 7.96
Attend-and-Excite - - - - 76.57 55.94 64.65 18.37 - 9.23
Divide-and-Bind - - - - 77.95 57.50 66.18 20.15 - 13.15
Our layout + Layout Guidance - - - - 71.09 45.98 55.84 16.58 - 29.85
Our layout + GLIGEN - - - - 75.34 62.14 69.01 23.72 - 48.41
Attention-Refocusing 88.62 91.91 90.44 77.05 77.43 61.78 68.72 25.60 89.57 48.29
DivCon (Ours) 94.17 89.53 91.80 80.73 78.65 63.41 70.22 29.97 91.87 54.21
Method
@@ 2.3 Text-to-Layout-to-Image Generation Models
As illustrated in Fig. 4, our framework comprises two stages. Given a text prompt, LLMs is first adopted to parse it and predict a layout. Then, the resultant layout is fed to the diffusion model as an additional condition with the text prompt to achieve text-to-image generation. More specifically, each stage is further divided into two steps to handle simple subtasks.
@@ 3.1 Text-to-Layout Prediction
Intuitively, to predict the layout from a text prompt, the number & spatial relations of the objects are first required to be reasoned, and then bounding boxes are generated to describe the coordinates and sizes of the objects. However, previous methods commonly conduct these two steps simultaneously without considering their inherent correlation, thereby hindering them from understanding complicated text prompts well. To remedy this, we employ the divide-and-conquer strategy to improve the accuracy of predicted layouts. Specifically, the text prompt is first fed to an LLM (e.g., GPT-4 (OpenAI 2024)) to perform numerical and spatial reasoning to produce more distinct instructions of objects, as illustrated in Fig. 4(a). Then, these instructions are adopted as additional conditions to predict more accurate bounding boxes.
(1) Numerical and Spatial Reasoning.
Given a text prompt $C$ describing an image, an LLM is adopted to predict a set of strings $O=\{s_{j}|j=1,2,...,n\}$ to describe the objects in the prompt. If the prompt primarily focuses on counting relationships among objects, each string $s_{j}$ includes both the name and quantity information of each object category $j$. If the prompt mainly describes spatial relationships, each string $s_{j}$ contains the object name and its spatial position in the image. After numerical and spatial reasoning, the numerical information and spatial relationships can be explicitly parsed from the complicated text prompt as intermedium conditions.
(2) Bounding Box Prediction.
After numerical and spatial reasoning, both the meta prompt $C$ and intermedium conditions $O$ are fed to the LLM again to predict a list of name-box pairs $P=\{p_{i}|i=1,2,...,n\}$. Each pair $p_{i}=\{o_{i}:(l_{i},t_{i},r_{i},b_{i})|i=1,2,...,n\}$ consists of the object name $o_{i}$ and its bounding box’s coordinates $(l_{i},t_{i},r_{i},b_{i})$. Meanwhile, an in-context exemplars including a process of correcting the features of generated bounding boxes is provided to the LLM to improve the quality of the predicted layouts.
@@ 3.2 Layout-to-Image Generation
After layout prediction, the resultant layout is employed as an additional condition to generate the image. As objects in the layout have different levels of difficulty to synthesize, a divide-and-conquer approach is adopted to reconstruct objects from easy ones to hard ones in an progressive manner, as shown in Fig. 4(b). First, the layout is fed to the diffusion model to synthesize all objects. Then, the consistency between the synthetic objects and the text prompt is calculated. Objects with low consistency are considered as low-fidelity ones and fed to the diffusion model for a second time while maintaining the results of objects with high consistency. With low-fidelity objects being highlighted, the diffusion model is encouraged to focus on these difficult objects in the second forward pass to achieve higher quality.
(1) First-Round Generation.
@@
We employ a layout-to-image diffusion model with sampling optimization (Phung, Ge, and Huang 2024) as the generative model to generate an image conditioned on the text prompt and the predicted layout. During the denoising process, random noise $z_{T}$ is fed to the diffusion model to gradually denoise it to an image $z_{0}$ after $T$ time steps.
(2) Consistency Evaluation.
@@
In the generated image $z_{0}$, each object is cropped using the bounding boxes in the layout and denoted as $x_{i}\in\mathbb{R}^{4\times h\times w}$, where $h\times w$ represents the height and width of the bounding box. Then, the CLIP similarity $S_{x_{i},o_{i}}\in(-1,1)$ is employed to evaluate the consistency between the cropped object and its name $o_{i}$ in the text prompt. The higher the similarity is, the higher the fidelity of the corresponding object is. Consequently, the objects with similarities higher than a specified threshold can be considered as easy samples that are ”good enough”. In contrast, other objects with low similarities are considered as hard samples that are more difficult to synthesize (e.g. ”sheep” and ”horse” in Fig. 3). Next, a binary mask $M$ is produced with bounding boxes of easy objects set to 1 and the rest regions set to 0.
(3) Second-Round Refinement.
@@
The objective of second-round refinement is to preserve the “good enough” objects while reproducing the remaining ones to ease the difficulty of the diffusion model. To this end, a straightforward approach is to retrieve the latent representation of the objects in the first-round denoising process. Nevertheless, directly freezing the object regions during the denoising process produces inharmonious artifacts around the boundaries. To address this issue, we retrieve the latent representation $z_{T}$ of all time steps during the denoising process of the first-round generation. Then, the binary mask $M$ is adopted to crop high-fidelity regions in the latent representation. At each time step $T$ of the second-round forward pass, these cropped representations are employed to replace the corresponding regions in the noises:
$z^{\prime}_{T}=z_{T}\circ M+z^{\prime}_{T}\circ(1-M)$ (1)
where $\circ$ is element-wise multiplication and $z^{\prime}_{T}$ is the noise at $T$ step of the second-round forward pass. With the latent representation of high-fidelity objects (masked by $M$) being retrieved, these objects can be well preserved to encourage the model to focus on reconstructing the remaining ones. Meanwhile, the bounding boxes of “bad samples” $P^{\prime}$ are passed to the diffusion model as the layout conditions, enabling the model to concentrate more on refining $P^{\prime}$ to further improve the overall quality of the generated image.
Methods Numerical Layout Pre. Rec. F1 Acc. Image Pre. Rec. F1 Acc. Spatial Layout Image Acc. Acc.
Stable Diffusion 2.1 - - - - 78.94 75.42 77.41 33.99 - 22.97
Attend-and-Excite - - - - 79.21 78.62 78.91 39.51 - 29.61
Divide-and-Bind - - - - 84.99 79.13 81.96 32.35 - 24.00
Our layout + Layout Guidance - - - - 81.97 75.41 78.55 43.13 - 52.88
Our layout + GLIGEN - - - - 85.13 84.97 85.05 52.21 - 59.72
Attention-Refocusing 98.51 99.11 98.81 94.31 84.61 85.64 85.12 56.10 95.05 69.28
DivCon (Ours) 98.47 98.35 98.41 94.39 85.41 87.06 86.22 57.35 96.23 73.15
Experiments
@@
In this section, we provide an evaluation of our method and compare it with state-of-the-art T2I models. Ablation experiments are then conducted to demonstrate the effectiveness of our divide-and-conquer approach in layout prediction and layout-to-image generation stages.